# TinyTim: A Family of Language Models for Divergent Generation

### v1

@@ Abstract
This work introduces TinyTim, a family of large language models fine-tuned on James Joyce’s ‘Finnegans Wake’. Through quantitative evaluation against baseline models, we demonstrate that TinyTim V1 produces a statistically distinct generative profile characterized by high lexical diversity and low semantic coherence. These findings are interpreted through theories of creativity and complex problem-solving, arguing that such specialized models can function as divergent knowledge sources within more extensive creative architectures, powering automated discovery mechanisms in diverse settings.
@@ 1 Introduction
Large Language Models (LLMs) based on the Transformer architecture (vaswani2017attention, ) have demonstrated powerful capabilities in synthesizing statistically likely patterns from vast data (brown2020language, ). However, this strength is also a fundamental limitation, constraining them to convergent, ‘mean-reverting’ outputs that inhibit the generation of genuinely novel hypotheses (dweck2023creativity, ). This issue reflects long-standing critiques of artificial reason, where formal systems struggle with the ambiguity and creative extrapolation inherent in human cognition (dreyfus1992, ; piantadosi2022meaning, ).
Human creativity is often described by dual-process theories that posit an interplay between focused, convergent thinking for evaluation and associative, divergent thinking for ideation (guilford1967nature, ; sternberg1985implicit, ). This cognitive duality is linked neuroscientifically to the dynamic coupling of the brain’s Executive Control Network (ECN) and Default Mode Network (DMN) (beaty2016creative, ; andrews201813, ; chen2025dynamic, ; LUCHINI2023195, ). While standard LLMs excel at tasks analogous to convergent thinking, they lack an intrinsic mechanism for the spontaneous, associative thought characteristic of divergence. To address this deficit, we engineered a divergent generative process into a language model by fine-tuning it exclusively on James Joyce’s ‘Finnegans Wake’—a text that is itself a product of extreme associative thought.
@@ 2.1 Data and Training
We fine-tuned the ‘TinyLlama-1.1B-Chat-v1.0’ model on the complete text of ‘Finnegans Wake’ (approx. 1.5MB). The text was preprocessed into 100-word segments to preserve its associative structure. The model, named TinyTim, was trained using a standard causal language modeling objective. TinyTim V1 has been available on HuggingFace for over a year 111hf.co/npc-worldwide/TinyTimV1, and has been downloaded more than 750 times in that time span.
In this work we showcase and characterize the V1 of the TinyTim family, and we plan to soon train and release an instruction-tuned version that can not only generate text like Joyce but can answer questions and be helpful as well to further facilitate creative processes.
@@ 2.2 Evaluation Framework
To quantify TinyTim’s generative profile, we compared it against a suite of baseline models trained for coherent responses: ‘qwen3:0.6b’, ‘llama3.2’, and ‘gpt-5-mini’. We generated responses from each model using a set of 10 creative prompts. Each response was evaluated using syntactic and semantic metrics:
@@ 3 Results
The generative profiles of TinyTim and the baseline models were evaluated across a suite of syntactic, lexical, and semantic metrics. The analysis reveals a statistically significant and functionally distinct separation between TinyTim and all conventionally trained models.
From an initial set of 2400 generated samples, 1013 were deemed valid for analysis after filtering out malformed or single-word outputs. The final sample sizes were: TinyTim (n=714), ‘gpt-5-mini’ (n=99), ‘llama3.2’ (n=100), and ‘qwen3:0.6b’ (n=1100). A Kruskal-Wallis test confirmed a significant difference ($p<.0001$) among the model groups for all metrics. Subsequent pairwise Mann-Whitney U tests with Bonferroni correction found that TinyTim was statistically different from all three baseline models on every metric.
Analysis of lexical sophistication metrics reveals a fundamental difference in how the models use vocabulary. As shown in Figure 1, while ‘gpt-5-mini’ possesses the largest raw vocabulary size, TinyTim demonstrates a radical degree of lexical invention.
Its Hapax Legomena Ratio, the proportion of words used only once, is over 50% higher than ‘gpt-5-mini’ (0.643 vs. 0.413), and its Yule’s K score—a robust measure of lexical richness that accounts for repetition—is over four times greater (208 vs. 47). This indicates TinyTim’s generative process is not based on retrieving a wide range of common words from a large vocabulary, but on constantly producing novel, rare, and often neologistic terms from a more constrained set of inputs. The baseline models are sophisticated retrievers; TinyTim is a lexical inventor.
@@
Figure 1: Measures of lexical sophistication. TinyTim is a clear outlier in lexical invention (Hapax Ratio, Yules K), distinguishing between creative generation and simple vocabulary size.
The distributional profiles of the models, shown in Figure 2, highlight the difference between convergent and divergent systems. The baseline models (‘gpt-5-mini’, ‘llama3.2’, ‘qwen3:0.6b’) produce outputs with tight, narrow, and predictable distributions for all metrics. This is the statistical signature of a reliable, convergent system optimized for consistency.
@@
In contrast, TinyTim’s distributions are extremely wide, heavily skewed, and possess long tails. This is most apparent in Unique Word Ratio, where its distribution spans a range nearly double that of any baseline, and in Sentence Complexity, where its outputs range from single-word fragments to sentences over 200 words long. This high variance is not a flaw, but the primary indicator of its function as an unpredictable, divergent generator whose outputs are intentionally varied and chaotic.
@@
Figure 2: Ridge plots showing the distribution of scores for the four primary metrics. The baseline models show tight, predictable distributions. TinyTim exhibits extreme variance and long tails, indicative of a divergent generative process. The x-axis represents the score for the corresponding metric.
The relationships between metrics further illuminate the different generative strategies. Figure 3 shows the correlations between key metrics across the full dataset. The most significant finding is in the top-left panel, which plots Token Diversity against Unique Word Ratio.
@@
The baseline models form a tight cluster in the high-diversity, low-uniqueness space. They draw from a vast, general vocabulary (high Token Diversity) but exhibit more repetition and less novelty within any single response (lower Unique Word Ratio). TinyTim occupies the opposite corner of the plot, forming a dispersed cloud in the high-uniqueness, low-diversity space. This is confirmed by the strong negative correlation for the full dataset (r=0.807r=-0.807). This illustrates the core paradox of its creativity: each individual response is highly novel (high UWR), but the overall pool of words it uses is highly specialized and constrained to its training data (low Token Diversity). This trade-off is the defining feature of its specialized, divergent function.
@@
Figure 3: Scatter plots showing relationships between primary metrics. The top-left panel highlights the fundamental trade-off between Token Diversity and Unique Word Ratio, clearly separating TinyTim’s divergent strategy from the convergent strategy of the baseline models.
@@ 5 Conclusion
This work has demonstrated that the generative properties of a language model can be fundamentally reshaped through targeted fine-tuning on radically unconventional literature. By analyzing the resultant model, TinyTim, against a suite of baselines, we have established a set of core findings regarding the nature of artificial creativity and model specialization. Our major conclusions are the following:
@@
    The generative profile of TinyTim establishes a formal distinction between the convergent, retrieval-based sophistication of standard LLMs and a divergent, combinatorial creativity that achieves novelty by exhaustively exploring the latent space of a focused, complex domain.
@@
    The model’s unique profile serves as a proof-of-concept for a specialized "divergent knowledge source," validating the architectural principle of using heterogeneous multi-agent systems for complex problem-solving.
  4. 4.
    The model’s function as a generator of high-variance, low-coherence semantic material implies a new paradigm for human-AI interaction, shifting from a query-response dynamic to a co-creative partnership where the user (or another convergent LLM) must perform the interpretive act.

### v2

@@ Abstract
In the search for artificial general intelligence, model development and training has focused primarily on vast datasets of known problems and their accepted solutions. This process necessarily produces convergent systems which are fundamentally incapable of the conceptual reframing that is required for genuine creative breakthroughs. Inspired by the divergent cognitive processes that allow humans to make such creative leaps, our work introduces a family of language models, TinyTim, to serve as sources of divergent generation within broader systems. These models have been created by fine-tuning on the anti-parsimonious text of James Joyce’s ‘Finnegans Wake’. Quantitative analysis of both an unsupervised fine-tuned model (TinyTim-V1) and a new instruction-tuned variant (TinyTim-V2) demonstrates a profound capacity for lexical invention; the foundational V1 model exhibits a Yule’s K score for lexical richness over twenty times greater than that of convergent baselines. This trait is a stable property of the family, as the instruction-tuned V2 maintains a statistically distinct profile and resists factual convergence, sacrificing benchmark performance to preserve its core generative style. This work establishes a methodology for engineering specialized divergent models that, when paired with convergent systems, can reframe problems and force breakthroughs beyond the reach of statistical optimization alone.
@@ 1 Introduction
Large Language Models (LLMs) based on the Transformer architecture (vaswani2017attention, ) have demonstrated powerful capabilities in synthesizing statistically likely patterns from vast data (brown2020language, ). However, this strength is also a fundamental limitation, constraining them to convergent, ‘mean-reverting’ outputs that inhibit the generation of genuinely novel hypotheses (dweck2023creativity, ). This issue reflects long-standing critiques of artificial reason, where formal systems struggle with the ambiguity and creative extrapolation inherent in human cognition (dreyfus1992, ; piantadosi2022meaning, ). Human creativity is often described by dual-process theories that posit an interplay between focused, convergent thinking for evaluation and associative, divergent thinking for ideation (guilford1967nature, ; sternberg1985implicit, ). This cognitive duality is linked neuroscientifically to the dynamic coupling of the brain’s Executive Control Network (ECN) and Default Mode Network (DMN) (beaty2016creative, ; andrews201813, ; chen2025dynamic, ; LUCHINI2023195, ). While standard LLMs excel at tasks analogous to convergent thinking, they lack an intrinsic mechanism for the spontaneous, associative thought characterisatic of divergence. To address this deficit, we first sought to isolate a divergent generative process by fine-tuning on James Joyce’s ‘Finnegans Wake’—a text that is itself a product of extreme associative thought. We then tested the capability of this process for conversational applications by integrating it into an instruction-tuned model, thereby creating and characterizing a family of specialized divergent tools.
@@ 2.1 Data and Training
The foundational model of the family, TinyTim V1, was created by fine-tuning the ‘TinyLlama-1.1B-Chat-v1.0’ model on the complete text of ‘Finnegans Wake’ (approx. 1.5MB). The text was preprocessed into 100-word segments to preserve its associative structure, and the model was trained using a standard causal language modeling objective. TinyTim V1 has been available on HuggingFace for over a year 111hf.co/npc-worldwide/TinyTimV1, and has been downloaded more than 750 times in that time span.
Building on this foundation, we then developed TinyTim-V2-IT  [^2] https://huggingface.co/npc-worldwide/tinytim-v2-1b-it to test if the divergent style could be harnessed within a conversational framework. This instruction-tuned variant, based on ‘google/gemma-3-1b-it’, was trained using TruthfulQA questions with answers rewritten in a Joycean style (100 epochs, learning rate 2e-4, LoRA rank 128), forcing the model to apply its divergent style to a convergent task.
In this work, we showcase and characterize both of these models.
@@ 2.2 Evaluation Framework
To quantify TinyTim’s generative profile, we compared it against a pair of baseline models trained for coherent responses: ‘qwen3:0.6b’ and ‘llama3.2’. We generated responses from each model using a set of 10 creative prompts. Each response was evaluated using syntactic and semantic metrics:
@@ 3 Results
The generative profiles of the TinyTim models and their baselines reveal a statistically significant and functionally distinct separation between divergent and convergent systems. From an initial set of generated samples, we analyzed a filtered dataset of 371 high-quality responses: TinyTim-V1 (n=73), TinyTim-V2-IT (n=100), ‘llama3.2’ (n=98), and ‘qwen3:0.6b’ (n=100). A Kruskal-Wallis test confirmed a significant difference ($p<.0001$) among the model groups for all metrics, with subsequent pairwise tests showing that both TinyTim models were statistically distinct from the baselines on every metric.
Analysis of lexical sophistication reveals the core of the divergent modification. As shown in Figure 1, while a model like ‘llama3.2‘ possesses a vast vocabulary (4084 unique words in our sample), the TinyTim family demonstrates radical lexical invention. The foundational model, TinyTim-V1, achieves a Yule’s K score—a robust measure of lexical richness—of 753.3, a value over twenty-four times greater than ‘llama3.2‘ (30.6). The instruction-tuned TinyTim-V2-IT, while more constrained, still registers a Yule’s K of 120.4, quadruple that of the baselines. This shows that the TinyTim models are not simply retrieving from a large vocabulary; they are lexical inventors.
@@
Figure 1: Measures of lexical sophistication. The TinyTim family, particularly V1, are clear outliers in lexical invention (Hapax Ratio, Yules K), distinguishing their creative generation from the large-vocabulary retrieval of baselines.
@@
The distributional profiles of the models, shown in Figure 2, highlight the difference between convergent and divergent systems. The baseline models (‘llama3.2’, ‘qwen3:0.6b’) produce outputs with tight, narrow, and predictable distributions for all metrics. This is the statistical signature of a reliable, convergent system optimized for consistency. In contrast, the distributions for both TinyTim-V1 and V2-IT are extremely wide and heavily skewed. This is most apparent in Unique Word Ratio, where V1’s distribution is nearly double that of any baseline, and in Sentence Complexity, where V2-IT produces outputs ranging from short fragments to sentences over 200 words long. This high variance serves as an indicator of their function as ‘unpredictable’ generators.
@@
Figure 2: Ridge plots showing the distribution of scores for four primary metrics. The baseline models show tight, predictable distributions. Both TinyTim models exhibit extreme variance and long tails, indicative of a divergent generative process. The x-axis represents the score for the corresponding metric.
@@
The relationships between metrics further illuminate the different generative strategies. Figure 3 shows the correlations between key metrics. The most significant finding is in the top-left panel, which plots Token Diversity against Unique Word Ratio. The baseline models form a tight cluster in the high-diversity, low-uniqueness space. They draw from a vast, general vocabulary (high Token Diversity) but exhibit more repetition within any single response (lower Unique Word Ratio). TinyTim-V1 occupies the opposite corner, forming a dispersed cloud in the high-uniqueness, low-diversity space, confirmed by the strong negative correlation for the full dataset (r=0.707r=-0.707). TinyTim-V2-IT aligns more with the baseline models in this view, but diverges from them in sentence complexity and average word length. These results illustrate the core paradox of their capabilities: each individual response is highly novel (high UWR), but the overall pool of words they use is highly specialized and constrained to their training data (low Token Diversity). This trade-off is the defining feature of their specialized, divergent function.
@@
Figure 3: Scatter plots showing relationships between primary metrics. The top-left panel highlights the fundamental trade-off between Token Diversity and Unique Word Ratio, clearly separating the divergent strategy of the TinyTim family from the convergent strategy of the baseline models.
3.1 Benchmark Performance as a Confirmation of Divergence
The stability of this divergent cognitive style was confirmed by testing TinyTim-V2-IT on a factual benchmark. It achieved only 52% accuracy on AI2 ARC-Easy, compared to baselines: ‘llama3.2’ (91%), ‘qwen3:0.6b’ (89%), and ‘gemma3:1b’ (80%). The model’s characteristic failure patterns included incoherent language, rambling responses, and hallucinated content, all evidence of its Joycean training overriding the instruction to be factual. This catastrophic failure is not a bug, but the critical piece of evidence proving that the divergent mode is a deep-seated, stable property of the model that resists convergence. It demonstrates a functional trade-off: the model sacrifices factual reliability to preserve its core generative style.
@@ 5 Conclusion
This work has demonstrated that the generative properties of a language model can be fundamentally reshaped through targeted fine-tuning on radically unconventional literature. By analyzing the TinyTim family of models, we have established a set of core findings regarding the engineering of specialized tools for augmenting current AI systems with more creative capabilities. Our major conclusions are the following:
@@
    The generative profile of the TinyTim family demonstrates their adherence to the uniqueness of their source material. Even in the instruction-tuned version, the characteristic style and inventiveness persist while functionally performing the tasks of responding to user prompts. Though its performance on the benchmark is degraded compared to its baseline ‘gemma3:1b’ model (76% versus 50%), it still performs better than the ‘gemma3:270m’ model (26%) and thus motivates further fine-tuning and training on larger local models like ‘gemma3:27b’ as well as the notion of a TinyTim-based reasoning model.
@@
    The models’ function as a generator of high-variance, low-coherence semantic material implies a new paradigm for human-AI interaction, shifting from a query-response dynamic to a co-creative partnership where the AI tool forces ‘out-of-the-box’ thinking. Specifically, the conversational capabilities of TinyTim-V2-IT demonstrate that this co-creative partnership can be engineered into AI systems that assist not by providing answers, but by forcing a deeper, more divergent engagement with the problem itself.